OpenStudy (wiggler):

Radius and height for smallest surface area given volume of an open cone is 27cm^3?

OpenStudy (dan815):

what is the surface area of a cone equaton

OpenStudy (wiggler):

For an open cone: \[SA = \pi r h\]

OpenStudy (dan815):

okay so lets optimize

OpenStudy (dan815):

what isthe volume of a cone formula?

OpenStudy (wiggler):

Volume: 27=13 basearea(r2) Base Area: A=πr2 Base area into Volume: 27=13πr2h So: h=81πr2 Putting that into Surface Area: SA=πr81πr2 =81r

OpenStudy (wiggler):

But taking the derivative of SA: SA′=−81r2 And since SA is at a min/max when SA' = 0: r2=−81 And then: r=9i Which doesn't work in the question.

OpenStudy (dan815):

that simplification is wrong

OpenStudy (wiggler):

Sorry, 13s should be 1/3, and the ^2 didn't work.

OpenStudy (wiggler):

Let me try again...

OpenStudy (dan815):

81 pi^2 * r^3

OpenStudy (wiggler):

Volume: \[27=\frac{ 1 }{ 3 } basearea(r^2)\] Base Area: \[A=πr^2\] Base area into Volume: \[27=\frac{ 1 }{ 3 }πr^2h\] So: \[h=\frac{ 81 }{ πr^2 }\] Putting that into Surface Area: \[SA=\frac{ πr81 }{ πr^2 }\] \[=\frac{ 81 }{ r }\]

OpenStudy (dan815):

ok

OpenStudy (wiggler):

So far so good?

OpenStudy (wiggler):

But taking the derivative of SA: \[SA′=\frac{ -81 }{ r^2 }\] And since SA is at a min/max when SA' = 0: \[r^2=−81\] And then: \[r=9i\] Which doesn't work in the question.

OpenStudy (dan815):

tha tsimplficiation is not right

OpenStudy (dan815):

u will get 1/r = 0 instead but even then

OpenStudy (wiggler):

At which step?

OpenStudy (dan815):

-81/r^2 = 0

OpenStudy (dan815):

dividing both sides and sqrting both side gives u 1/r = 0

OpenStudy (wiggler):

okay, so what should I do?

OpenStudy (dan815):

well this still isnt giving u very nice answers let me try doing it from start see if there is some problem

OpenStudy (dan815):

because right now if we solve this.. r= +/- inf and h will be going to 0 which doesnt make sense

OpenStudy (dan815):

i see ur problem

OpenStudy (dan815):

the SA =pi r*(sqrt(h^2+r^2))

OpenStudy (wiggler):

But that's for a cone including the base, isn't it? The question asks for a "cone shaped paper drinking cup" so shouldn't it be open? or does this not matter?

OpenStudy (dan815):

oh u are right

OpenStudy (wiggler):

Okay, no, that is the equation for an open cone. Right, I'll try again...

OpenStudy (dan815):

oh okay yeah

OpenStudy (dan815):

becayse it SA= pi*r*L

OpenStudy (dan815):

L=sqrth^2+r^2

OpenStudy (wiggler):

Yep :)

OpenStudy (wiggler):

Surface Area (given open top): \[SA=πr \sqrt{h^2+r^2} \] Volume: \[27=\frac{ 1 }{ 3 } basearea(r^2)\] Base Area: \[A=πr^2\] Base area into Volume: \[27=\frac{ 1 }{ 3 }πr^2h\] So: \[h=\frac{ 81 }{ πr^2 }\] Putting that into Surface Area: \[SA=πr \sqrt{\frac{ 81 }{ \pi r^2 }+r^2}\]

OpenStudy (dan815):

ok

OpenStudy (dan815):

bring the pi r into sqrt and simplify

OpenStudy (wiggler):

\[SA = \sqrt{(\frac{ 81 }{ \pi r^2 }+r^2)\pi^2 r^2}\] Is that right?

OpenStudy (dan815):

yep simplify that now

OpenStudy (dan815):

d/dr sqrt(pi 81 +pi^2 r^3) = 3pi^2r^2/2 (pi 81 +pi^2 r^3)^(-1/2)

OpenStudy (wiggler):

Shouldn't it be d/dr sqrt(pi 81 + pi^2 r^((4)))? (changed 3 to 4 in extra brackets)

OpenStudy (wiggler):

So: \[SA = \sqrt{81\pi + r^4 \pi^2}\] \[SA' = \frac{ 1 }{ 2 }(81\pi + r^4 \pi^2)^{-1/2}(4r^3 \pi^2)\]

OpenStudy (wiggler):

But then SA' = 0 still gives r=0

OpenStudy (wiggler):

What went wrong?

OpenStudy (anonymous):

what formula did u use for SA?

OpenStudy (wolf1728):

See 3rd post down PI*r*h

OpenStudy (anonymous):

that's incorrect

OpenStudy (wiggler):

No, changed later to SA = pi r sqrt(h^2 + r^2)

OpenStudy (wolf1728):

Sorry wiggler

OpenStudy (wiggler):

Haha, all good :) It's a LONG work in progress

OpenStudy (anonymous):

did u substitute h or r above with the volume?

OpenStudy (wiggler):

substituted in h = 81/r^2 to the SA eqn

OpenStudy (wiggler):

(eliminated h in order to find r)

OpenStudy (anonymous):

u missing the pi in ur substiute, u sure of it?

OpenStudy (wolf1728):

well gotta go - seems kx2bay will be a BIG help. see you folks :-)

OpenStudy (wiggler):

Where is the pi missing?

OpenStudy (anonymous):

u said above "substituted in h = 81/r^2 to the SA eqn", where is the pi ? volume = 27 = 1/3 pi r2 h

OpenStudy (wiggler):

Oh, yeah, I did include that in the calculations.

OpenStudy (wiggler):

Copying error

OpenStudy (wiggler):

That gave me SA = pi r * sqrt((81/pi r^2) + r^2)

OpenStudy (anonymous):

I started with subst. r = 9/sqrt(pi.h) and got at the end this \[9 (81.h ^{-2} + \pi.h)^{\frac{ 1 }{ 2 }}\] this is where am struggling to find the derivative, do u know what it will be?

OpenStudy (wiggler):

\[4.5(81h^-2 + \pi h)(-162h^-3 + \pi)\]

OpenStudy (wiggler):

That should be h^-2 and h^-3

OpenStudy (anonymous):

thanks first answer for h is negative i.e. h = cubeRoot(-pi/81)

OpenStudy (anonymous):

second answer for h is good i.e. h^3 = 162/pi

OpenStudy (anonymous):

so h = 3.722 cm what is r ? based on r=9/sqrt(pi.h)

OpenStudy (anonymous):

i got r = 2.632 cm lets check the volume now if it works

OpenStudy (anonymous):

yep I get 27 for the volume so that's ur answers h = 3.722 and r = 2.632 now teach me how did you work out the derivative above that got stuck on because I need to refresh my rusty memory on that :) please...

OpenStudy (anonymous):

that I got stuck on

OpenStudy (wiggler):

Sure :) You start with the outer function (the power of 1/2) Taking the stuff in the brackets as "x", take the derivative of 9x^(1/2)

OpenStudy (anonymous):

ok, continue please

OpenStudy (anonymous):

it will give u 1/2 *9*x^-1/2

OpenStudy (wiggler):

Yep. Then multiply that by the derivative of the inner term. i.e. what you just calculated * derivative of (81/h^2 + pi h)

OpenStudy (anonymous):

hmmm u too good with stuff, thanks for the tutoring :)

OpenStudy (wiggler):

No problem :) just remember to start with the outermost function and work your way in. And thanks for your help!

OpenStudy (anonymous):

no worries and thank you cheers